OpenStudy (anonymous):

Find the zeroes of the function algebraically f(x)=x^2+2x-8 Please show step by step!

zepdrix (zepdrix):

We have roots (places where the function crosses the x-axis) when f(x)=0. That will be our first step, let f(x)=0.\[\large 0=x^2+2x-8\]

zepdrix (zepdrix):

Are you comfortable with factoring? This quadratic can be factored quite nicely, it takes a little bit of practice to recognize the right numbers though. :D

OpenStudy (anonymous):

factoring? I'm not quite sure what you mean.. It sounds familiar. Can you give me an example?

zepdrix (zepdrix):

Oh, they used the word "zeroes", im sorry i misread that. Zeroes and roots are the same thing.

zepdrix (zepdrix):

Sure :)

zepdrix (zepdrix):

\[\large 0=x^2+5x+6\] The idea is, we want to be able to break this down into 2 factors of the form:\[\large 0=(x \pm a)(x \pm b)\] You want FACTORS of the 6, that ADD to give you 5. Hmm, So we can get 6 by multiplying 1 and 6. But 1 and 6 add to 7, hmm that's no good. How about 3 times 2? Hmm, 3 and 2 add to make 5! Ooo I think that will work. Let's set it up,\[\large 0=(x+3)(x+2)\] If you wanted to verify, you could multiply these out just to check. Remember how to FOIL things? :)

OpenStudy (anonymous):

Oh! I remember this! But, I don't know what FOIL is

zepdrix (zepdrix):

It's an acronym telling you how to multiply 2 binomials together, (But I can't for the life of me remember what it stands for, lol).

OpenStudy (anonymous):

So what you would do for my problem is find what adds to 2 and multiplies to -8 x^2+2x−8 Sooo 4 and -2?

OpenStudy (anonymous):

And the equations would be (X+4) (X-2)

OpenStudy (anonymous):

right?

zepdrix (zepdrix):

Yesss very good \:D/

OpenStudy (anonymous):

So the zeroes if the function would be -4 and 2?

zepdrix (zepdrix):

Yes good, x=-4, x=2! :)

OpenStudy (anonymous):

Sweet! Thanks!